# SFO: Piloting VLM Feedback for Offline RL

### v4

@@ 5 Conclusion
Our study provides key insights into the integration of VLM feedback with offline RL. First, sub-trajectory selection is critical, as full-trajectory preference learning exacerbates the stitching problem, making effective credit assignment impossible. Using sub-trajectories allows for improved learning stability and the ability to extract high-quality behaviors from mixed datasets. Second, non-Markovian feedback is essential, even in Markovian environments, since VLMs lack an understanding of control dynamics. Instead, they require vision-based trajectory analysis to correctly assess improvement over time. Our results confirm that non-Markovian prompts lead to significantly better performance. Finally, SFBC consistently outperforms competitive baselines, demonstrating that treating VLM-derived success probabilities as weights rather than rewards leads to superior offline policy learning. Our retrospective filtering mechanism further enhances performance by removing failure-propagating sub-trajectories.
Our findings suggest that VLM feedback provides non-myopic guidance, mitigating the need for explicit reward propagation and addressing key challenges in scaling offline RL. Most immediately, this work could be applied to automate feedback across a sufficiently diverse set of tasks to support training an RL foundation model from offline data. Such an application of SFBC would inherently fall under the category of offline meta-RL (Beck et al., 2023), while also obviating the need for complex meta-gradient computation (Vuorio et al., 2021). Putting such guidance systems on autopilot may finally provide the runway needed to arrive at a truly generalist agent for RL.

### v5

# Offline RLAIF: Piloting VLM Feedback for RL via SFO

@@ 5 Conclusion
Our study provides key insights into the integration of VLM feedback with offline RL. First, sub-trajectory selection is critical, as full-trajectory preference learning exacerbates the stitching problem, making effective credit assignment impossible. Using sub-trajectories allows for improved learning stability and the ability to extract high-quality behaviors from mixed datasets. Second, non-Markovian feedback is essential, even in Markovian environments, since VLMs lack an understanding of control dynamics. Instead, they require vision-based trajectory analysis to correctly assess improvement over time. Our results confirm that non-Markovian prompts lead to significantly better performance. Finally, SFBC consistently outperforms competitive baselines, demonstrating that treating VLM-derived success probabilities as weights rather than rewards leads to superior offline policy learning. Our retrospective filtering mechanism further enhances performance by removing failure-propagating sub-trajectories. Our findings suggest that VLM feedback provides non-myopic guidance, mitigating the need for explicit reward propagation and addressing key challenges in scaling offline RL.
Most immediately, this work could be applied, at scale, to automate feedback across a sufficiently diverse set of tasks to support training a general RL foundation model from offline data. Training the RL agent to be adaptive would inherently fall under the category of offline meta-RL (Beck et al., 2023; 2025). Moreover, implementing the agent with a general-purpose sequence model, such as a transformer (Vaswani et al., 2017), would also place it within the domain of in-context RL (Moeini et al., 2025), while also obviating the need for complex meta-gradient computation (Vuorio et al., 2021). Putting feedback and guidance systems on autopilot may finally provide the runway needed to arrive at a truly generalist agent for RL.